OpenStudy (anonymous):

a telephone pole casts a shadow that is 9m long, at the same time a person that is 1.5 tall cast a shadow that is 0.6 m long. how tall in meters is the telephone pole

OpenStudy (anonymous):

This can be solved using ratios. \[\frac{ 9m }{ ? } = \frac{ 0.6m }{ 1.5m }\]

OpenStudy (anonymous):

if you still need help let me know

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

would it be 22.5?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

thank you so much , can you help me with another question ?

OpenStudy (anonymous):

i can try

OpenStudy (anonymous):

ok. when plane a is 60 miles south and 80 miles west of the airport and plane b is 30 miles north and 40 miles east of the airport how far away are the planes

OpenStudy (anonymous):



OpenStudy (anonymous):

Find the Hypotenuse for each triangle and add them together. Again you can see that they are similar triangles.

OpenStudy (anonymous):

A to B would be a straight line, even though the drawing doesn't look like it is.

OpenStudy (anonymous):

Do you need more help with this question?

OpenStudy (anonymous):

yes. im sorry im so confused.

OpenStudy (anonymous):

would it be 150 miles apart?

OpenStudy (anonymous):

ok. plane A is \[\sqrt{60^{2}+80^{2}}\] miles from the airport

OpenStudy (anonymous):

that is correct. 150.

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

you are welcome